OpenStudy (katherinesmith):

need help with solving linear equations with 3 variables

OpenStudy (katherinesmith):

Solve x + 10y – 2z = 20 5x – 3y + 4z = 8 2x + y = 6

OpenStudy (katherinesmith):

@phi i don't think this video is talking about the same thing i'm trying to figure out.

OpenStudy (katherinesmith):

no wait nevermind. i think i got it

OpenStudy (phi):

It is more advanced than what you have to do... but the idea is the same you start with x + 10y – 2z = 20 5x – 3y + 4z = 8 2x + y = 6 and write it as 1 10 -2 20 5 -3 4 8 2 1 0 6

OpenStudy (katherinesmith):

and then find the inverse and then multiply the inverse by 20, 8, 6. correct?

OpenStudy (phi):

yes, that works, but you only find the inverse of 1 10 -2 5 -3 4 2 1 0

OpenStudy (phi):

Finding the inverse of a 3 x 3 is a pain if you are doing it by hand

OpenStudy (phi):

was going to go through using elimination but the numbers are very ugly. though the final answers for x,y and z are simple

OpenStudy (phi):

The good thing about elimination is it is a formula you follow. The bad thing is you can get ugly numbers. But here are the gory details. 1 10 -2 20 5 -3 4 8 2 1 0 6 step 1: zero out the 5 in the 2nd row. multiply the first row by -5: -5 -50 10 -100 (notice the big numbers, ugh!) add this to the 2nd row to get : 0 -53 14 -92 1 10 -2 20 0 -53 14 -92 2 1 0 6 step 2: zero out the 2 in the 3rd row. multiply the first row by -2: -2 -20 4 -40 add this to the 3rd row to get: 0 -19 4 -34 1 10 -2 20 0 -53 14 -92 0 -19 4 -34 those are ugly numbers.

OpenStudy (phi):

1 10 -2 20 0 -53 14 -92 0 -19 4 -34 next, move to the 2nd column, 2nd row. step 3: make the -53 =1. We do this by dividing that row by -53. we get 1 10 -2 20 0 1 -14/53 92/53 0 -19 4 -34 step 4 zero out the -19 in the 3rd row. to do this, multiply the SECOND row by 19 to get 19 -14*19/53 19*92/53 (I don't want to use decimals and lose precision, but you can do so with a calculator) add that to the last row to get 1 10 -2 20 0 1 -14/53 92/53 0 0 -54/53 -54/53 step 5: make the -54/53 in the 3rd row a 1. we do this by multiplying that row by -53/54. we get 1 10 -2 20 0 1 -14/53 92/53 0 0 1 1 the last row looks nice. It means z = 1 we could continue with elimination, but we can solve the problem using *back substitution"

OpenStudy (phi):

the 2nd row 0 1 -14/53 92/53 means y - 14/53 z = 92/53 but we know z=1, so this is y - 14/53 = 92/53 solve for y= 92/53 + 14/53 = 106/53 =2 (a miracle, it came out so nice)

OpenStudy (phi):

now we move to the first row 1 10 -2 20 which means 1x + 10y -2z = 20 but we know y=2 and z= 1 so this is x +20 -2 = 20 x = 2 the final answer is x=2, y=2, z=1